OpenStudy (anonymous):

I'm just learning about using scientific notation to express base units. Can someone please walk me through this problem? 114 Pg

OpenStudy (abhisar):

Hello @Alyssa8 ! What's ur question ?

OpenStudy (anonymous):

Hello. I'm not entirely familiar with how to convert 114 Pg into scientific notation,

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

well what is 114 Pg ?

OpenStudy (anonymous):

is it Picograms?

OpenStudy (abhisar):

oh i see...!

OpenStudy (abhisar):

So, 114 Pg = 114 \(\sf \times 10^{-12}g\)

OpenStudy (anonymous):

Yeah, I'm pretty sure I have to convert it into grams using scientific notation. I'm just unsure of the process in which I get the answer.

OpenStudy (anonymous):

How did you get that answer, if I may ask?

OpenStudy (abhisar):

Now let's convert it into scientific notations, \(\sf 114 \times 10^{-12}=1.14 \times 10^2 \times 10^{-12}=1.14 \times 10^{-10}g\)

OpenStudy (abhisar):

The correct scientific notation is \(\sf 1.14\times 10^{-10}g\)

OpenStudy (anonymous):

Is the 10^-12 normal for units measure in picos?

OpenStudy (abhisar):

yes, 1Pg=\(\sf 10^{-12}g\)

OpenStudy (anonymous):

Okay, thank you. that makes much more sense.

OpenStudy (abhisar):

\(\rlap{\color{blue}{\huge\bigstar}\huge \color{blue}{ \text{You're Most Welcome! }}\color{blue}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{You're Most Welcome! }}\color{aqua}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (unklerhaukus):

if its a capital P it means peta- i.e. x10^15 if its a lowercase p it means picto- i.e. x10^-12

OpenStudy (unklerhaukus):

so your answers off by a factor of 10^27

OpenStudy (unklerhaukus):

method is right though.